OpenStudy (anonymous):

Multiply out the vertex form (y= a(x-h)^2 + k). Compare the h and k to the a, b, and c in the standard form (y=ax^2 +bx+c). What information do the parameters or combinations of parameters provide about the graph of the quadratic function?

OpenStudy (anonymous):

I am desperate and really need help! Would appreciate any help given.

OpenStudy (joannablackwelder):

Multiply everything out and try to make the equation look as much like ax^2 +bx + c as possible.

OpenStudy (joannablackwelder):

a(x-h)(x-h) +k

OpenStudy (joannablackwelder):

a(x^2 - 2xh + h^2) + k

OpenStudy (joannablackwelder):

Make sense so far?

OpenStudy (anonymous):

Yes. I got the same thing today. Now all I had to compare was standard form to what we got. So c=ah^2+k ax^2=ax^2 h+k= a+b+c. The question says to compare h and k to the a,b and c in standard form. IT also says what information does the parameters give us about the quadratic function. Is there anything else that I can add into my answer?

OpenStudy (joannablackwelder):

I got that a=a, b=-2ah, and c=ah^2+k

OpenStudy (joannablackwelder):

So, since the width of the parabola is determined by the leading coefficient of standard form, the h and k do not affect that term and thus don't affect the width of the graph.

OpenStudy (anonymous):

Is that the final answer?

OpenStudy (joannablackwelder):

That is the only thing I can think of. Not sure if your teacher is looking for something else...

OpenStudy (joannablackwelder):

@abb0t Any thoughts?

OpenStudy (abb0t):

Did you foil it out?

OpenStudy (anonymous):

That Answers my question thanks for the help!